If Amy spent her entire pocket money, she could afford 8 candy bars and 8 comic books a week. She could also just afford 10 candy bars and 4 comic books a week. The price of a candy bar is ₹ 5. What is Amy's weekly pocket money?






1
Expert's answer
2022-03-11T08:40:35-0500

"W=8br+8bk,"

"W=10br+4bk,"

"8br+8bk=10br+4bk,"

"2bk=br,"

"br=5,"

"bk=0.5br=2.5,"

"W=8\cdot 5+8\cdot 2.5=60."